OpenStudy (anonymous):

A ball is thrown into the air with an initial upward velocity of 45 ft/s. The height (h) in feet is given by the equation below. After about how many seconds will the ball hit the ground?

OpenStudy (anonymous):

h=-16t^2+45t+6

Nnesha (nnesha):

t represent time h = height what would be the height when ball hit the ground ?

Nnesha (nnesha):

 height = ?

OpenStudy (anonymous):

t?

Nnesha (nnesha):

t = time we need to find time but when ball hits the ground h =0 so replace h by 0 \[\huge\rm 0=-16t^2+45t+6\] and now solve for t you can use quadratic formula \[ \huge\rm \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

Nnesha (nnesha):

a b c Ax^2 +Bx +C = 0